A company tested 735 of the lightbulbs they produced and found them to have a mean life of 1,200 hours and a standard deviation of 50 hours. How many of these lightbulbs had a life between 1,170 hours and 1,230 hours?


1
Expert's answer
2021-01-18T06:12:51-0500


We assume that the distribution of the values of life of the lightbulbs can be well approximated by a normal distribution with the same parameters: mean 1,200 hours and standard deviation 50 hours.

Then the value of life is equal to 1200 + 50X, where X is a standard normally distributed random variable. The probability that the value of life is between 1,170 and 1,230 hours is equal to the probability that the value of X is between -0.6 and 0.6:

"1170 \leq 1200 + 50X \leq 1230"

"(1170-1200)/50 \leq X \leq (1230-1200)/50"

"-0.6\leq X \leq 0.6"

This probability may be found as the integral

"\frac{1}{\sqrt{2\pi}}\int\limits_{-0.6}^{0.6}e^{-x^2/2}dx = \Phi(0.6) -\Phi(-0.6) =0.72575 - 0.27425 = 0.4515"

where "\Phi(x) = \frac{1}{\sqrt{2\pi}}\int\limits_{-\infty}^{x}e^{-t^2/2}dt" - the cumulative distribution function of the standard normal distribution, its values we may get from the tables.

The estimated number of lightbulbs had a life between 1,170 hours and 1,230 hours is equal to "735 \cdot 0.4515 = 331.85"

Answer. The estimated number of lightbulbs had a life between 1,170 hours and 1,230 hours is equal to 331.85.